Project A requires an initial outlay at t = 0 of $4,000, and its cash flows are the same in Years 1 through 10. Its IRR is 16%, and its WACC is 11%. What is the project's MIRR? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Investment = 4000

Let equal annual cash inflow = X

Present value of these cash inflows at IRR must be equal to investment

X * [ 1 - (1.16)-10 ] / 0.16 = 4000

X * 4.83322748 = 4000

X = 4000 / 4.83322748

X = 827.60

Step - (2)

Let MIRR = X

Future value of current investment after 10 years = 4000 * ( 1 + x )10

Future value of cash inflows = 827.60 * [ (1.11)10 - 1 ] / 0.11 ...... Here r = WACC = 0.11  

= 13839.13

Now .............. 4000 * ( 1 + x )10 = 13839.13

( 1 + x )10 = 13839.13 / 4000 = 3.4597825

1 + x = 10th root ( 3.4597825 )

x = 1.1322

x = 13.22 %